OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 50 km/h. After six hours, the velocity of the car is 54 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points)

OpenStudy (anonymous):

ill give a medal i really need to finish thanks

OpenStudy (anonymous):

HELP PLEASE IVE BEEN WORKING SINCE 8 IN THE MORING

OpenStudy (anonymous):

find the slope for points (2, 50) and (6, 54)

OpenStudy (anonymous):

i dont care if its rigtht anymore i jsut need an answer

OpenStudy (anonymous):

please can you give an equation

OpenStudy (anonymous):

im super tired

OpenStudy (anonymous):

use slope formula (54-50)/(6-2) so m=1

OpenStudy (anonymous):

then use point slope form : y-50=x-2 so the answer should be x-y=-48

OpenStudy (anonymous):

I hope its right im not 100% sure though lol

OpenStudy (anonymous):

thanks so much i hope someone answered your question